OpenStudy (anonymous):

Triangles ABC and LMN are given such that AB = 24, BC = 20, and AC = 40 LM = 6, MN = 5, and LN = (3a + 1) To be similar, which should be the value of a? Answer 6 5 4 3

OpenStudy (anonymous):

a - 3.

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

just realise that 40/(3a+1)=4

OpenStudy (anonymous):

Notice thar LM is a quarter of AB, and MN is a quarter of BC. For the triangles to be similar, LN must be a quarter of AC, or 3a + 1 = 40/4.